OpenStudy (anonymous):

perform the indicated operation; simplify. 5p^5q^2/3p^4 divided by 15p/q^6

OpenStudy (anonymous):

is this \[\frac{5p^5q^2}{3p^4}\div \frac{15p}{q^6}\]

OpenStudy (anonymous):

if so, invert and multiply but cancel first. \[\frac{5p^5q^2}{3p^4} \times \frac{q^6}{15p}\] \[=\frac{5p^5q^8}{30p^5}\] \[=\frac{q^8}{6}\]

OpenStudy (anonymous):

thankyou

OpenStudy (gw2011):

I believe the answer is q^8/9

OpenStudy (anonymous):

oops i am wrong and you are right. the denominator is 9 not 6. sorry.